OpenStudy (anonymous):

Help with a word problem please. "During a season, a football team scored a total of 168 points on touchdowns, extra points, and field goals. A touchdown is worth 6 points, an extra point is worth 1, and a field goal is worth 3. The team had a total of 48 scoring plays and scored twice as many touchdowns as field goals. Write a system to represent the teams scoring during the season, let x=touchdowns, y= extra points, and z=field goals." I need help getting the system for this problem then I can solve the rest, thanks :3

OpenStudy (anonymous):

Points from touchdowns + Pts from extra pts + Pts from field goals = 168

OpenStudy (anonymous):

What's an expression for "Points from touchdowns"?

OpenStudy (anonymous):

Points from touchdowns....that's just 6x points from extra pts and points from field goals give us similar expressions. So what equation can we write _______ + ______ + ________ = 168

OpenStudy (texaschic101):

6x + y + 3z = 168 x = 2z I am just not sure what to do with the 48 plays

OpenStudy (anonymous):

Okay I get that for the first equation, but I need a system, what other 2 equations can I find in that word problem? I'm not sure what to do with the 48 either

OpenStudy (anonymous):

x + y + z = 48

OpenStudy (texaschic101):

oh....I get it now

OpenStudy (anonymous):

no coefficients?

OpenStudy (anonymous):

6x + y + 3z = 168 x = 2z x + y + z = 48

OpenStudy (anonymous):

Do yo understand where the equations come from?

OpenStudy (anonymous):

Yeah I do, thank you. I was confused from the word problem. Thanks guys for the help :D

OpenStudy (texaschic101):

give BangkokGarrett a medal....he deserves it